princeevee:

A stock bottle of concentrated hydrochloric acid is a 37% (w/w) HCl solution. What is the molarity of this solution? How many moles of HCl are in a 10 mL sample of this concentrated acid? The density of 37% HCl is 1.19 g/mL.

princeevee:

@563blackghost

Nnesha:

Hi.

Nnesha:

\[\large\rm Density = \frac{mass}{volume}~~~,~~~~~ Molarity =\frac{moles~of~solute}{Liters~of~solution}\] first we have to find moles of HCL given that there are 37 g HCL in 100 g of solution

Nnesha:

are these two separate questions or one ?

Nnesha:

1 mol of HCL = 36.46094 g given that 37% is the HCL solution \[\large\rm 37 g-HCl \cdot \frac{1 ~mol ~of~HCl}{36.46094 ~g ~of~HCl}\] \[\rm 37 g-HCl \cdot \frac{1 ~mol ~of~HCl}{36.46094 ~g ~of~HCl}=1.01478\] lowest decimal place is 2 so it will be 1.01 now use the equation density=mass/volume to find the volume \[\large\rm 1.19 \frac{g}{mL} = \frac{100~g}{~~volume }\] solve for volume \[\rm volume =\frac{100~g}{ 1.19~\frac{g}{mL}}= \frac{100~g}{1.19~g}\cdot mL\] \[\rm volume = = \frac{100~\cancel{g}}{1.19~\cancel{g}}\cdot mL=84.03~ mL\] convert ml to liter so the volume =0.084 Liter Use the morality equation \[\large\rm \color{red}{Molarity }= \frac{moles ~of~solute}{Liters~of~solution}=\frac{1.01~g}{0.084 ~L}\]

Nnesha:

once you calculate that ^^molarity use the same equation to find moles of solute